Data are given below on the adjusted gross income x and the amount of itemized deductions taken by taxpayers. Data were reported in thousands of dollars. With the estimated regression equation = 4.68 + .16x, the point estimate of a reasonable level of total itemized deductions for a taxpayer with an adjusted gross income of $52.5 thousand is $13.08 thousand.

Adjusted Gross Income($1000s) Reasonable Amount of Itemized Deductions ($1000s)
22 9.6
27 9.6
32 10.1
48 11.1
65 13.5
85 17.7
120 25.5

In the questions that follow, enter the dollar amounts requested. For example, if the regression results provide a value of 11.74 thousand, enter 11,740 as the dollar amount in the box. Use Table 1 of Appendix B.

a)Develop a 95% confidence interval for the amount of total itemized deductions for all taxpayers with an adjusted gross income of $52,500 (to the nearest whole number).

$ ( ____ , _____ )

b)Develop a 95% prediction interval for the amount of total itemized deductions for a particular taxpayer with an adjusted gross income of $52,500 (to the nearest whole number).

$ ( ____ , _____ )

c)If the particular taxpayer referred to in part (b) claimed total itemized deductions of $20,400, would the IRS agent's request for an audit appear to be justified?

d)Use your answer to part (b) to give the IRS agent a guideline as to the amount of itemized deductions that would suggest an audit for a taxpayer with an adjusted gross income of $52,500. Any deductions exceeding the upper limit of $_______ could suggest an audit.
Answer #1


Incomex Deductions x2 SUMMARY OUTPUTi 9.6 9.6 10.1 11.1 13.5 17.7 25.5 27 32 48 65 85 729 1024 2304 4225 7225 14400 30391 Reg

y=Deductions;~x=Income\\ \hat{y}=4.6768+0.1613x\\ \bar{x}=46.5,~S_{xx}=\sum x_i^2-7\bar{x}^2=30391-7*46.5^2=15255.25\\ t_{0.025,5}=2.5706\\ At~x=52.5~then~\hat{y}=4.6768+0.1613*52.5=13.1451,\\~s=standard~error=1.3716\\ \\ (a)~95\%~C.I.~for~the~amount~of~total~itemized~deductions~for~all~tax~payers~with\\ adjusted~gross~income~of~\$52500~is\\ \left((\hat{y}-t_{0.025,5}\times s\times \sqrt{\frac{1}{7}+\frac{(52.5-\bar{x})^2}{S_{xx}}})*1000,~\hat{y}+t_{0.025,5}\times s\times \sqrt{\frac{1}{7}+\frac{(52.5-\bar{x})^2}{S_{xx}}})*1000 \right )\\ =(\$11802,~\$14489)\\ (b)~95\%~P.I.~for~the~amount~of~total~itemized~deductions~for~all~tax~payers~with\\ adjusted~gross~income~of~\$52500~is\\ \left((\hat{y}-t_{0.025,5}\times s\times \sqrt{1+\frac{1}{7}+\frac{(52.5-\bar{x})^2}{S_{xx}}})*1000,~\hat{y}+t_{0.025,5}\times s\times \sqrt{1+\frac{1}{7}+\frac{(52.5-\bar{x})^2}{S_{xx}}})*1000 \right )\\ =(\$9372,~\$16918)\\

(c) Since total itemized deductions of $20,400 aboce the 95% P.I. so the IRS agent's would request for an audit appear to be justified.

(d) Use your answer to part (b) to give the IRS agent a guideline as to the amount of itemized deductions that would suggest an audit for a taxpayer with an adjusted gross income of $52,500. Any deductions exceeding the upper limit of $16918 could suggest an audit.